Find the sample size needed to estimate the percentage of democrats among registered voters in Florida. Use a 0.01 margin of error, and use a confidence level of 96% and assume p=0.55
Answer #1

given data are:-

p = 0.55

margin of error (E) = 0.01

z critical value for 96% confidence level, both tailed test be:-

z^*=2.054[ in any blank cell of excel type =NORMSINV(0.98)]

the needed sample size be:-

=\left ( \frac{z^*}{E} \right )^2*p*(1-p)

=\left ( \frac{2.054}{0.01} \right )^2*0.55*(1-0.55)

\approx\mathbf{10442}

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...